# Fluid dynamics, Statistical physics, Theoretical physics

### v2

@@ Acknowledgements
I am grateful for insightful discussions regarding this theory with Herman Verlinde, who offered several inspiring remarks. Conversations with Katepalli Sreenivasan concerning the physics of turbulence, as observed in direct numerical simulations and laboratory experiments, were invaluable in clarifying the role and significance of loop space diffusion.
@@ Appendix A Essentials of the loop space calculus
The full derivations and technical details can be found in our recent paper [Migdal(2025b)].
@@
The area derivative is defined by the non-abelian Stokes theorem. For a detailed derivation, see [Migdal(2025b)], Sec. 3.3.
@@ A.1 Operator representation of the Wilson loop
The loop functional can be represented as a trace of a path-ordered exponential of covariant derivative operator. This subsection proves the fundamental operator identity that enables this. The full derivation, originally presented in [Migdal(2025b)] (Sec. 3.4), is reproduced here with all necessary mathematical details. We also fix some typos in that paper.
@@
The proof relies on discretizing the path and using the "disentangling" identity, which is a consequence of Feynman’s operator calculus [Feynman(1951)].
@@ 4. The Closed Loop Condition
The field strength $F_{\mu\nu}$ is generated by the antisymmetric part of the second dot derivative. For details, see [Migdal(2025b)], Sec. 3.6, Eq. (25).
@@
The covariant derivative of the field strength is generated by the discontinuity of the dot derivative, $\partial_{\mu}(t)\equiv\frac{\delta}{\delta\dot{C}_{\mu}(t-)}-\frac{\delta}{\delta\dot{C}_{\mu}(t+)}$. For details, see [Migdal(2025b)], Sec. 3.7, Eq. (31).
@@
The dot derivative discontinuity $\partial_{\mu}(t)$ annihilates the Wilson loop. For details, see [Migdal(2025b)], Sec. 3.9, Eq. (33).
@@
The area derivative acts as a first-order differential operator, satisfying the standard Leibniz rules. For details, see [Migdal(2025b)], Sec. 3.10, Eqs. (39-41).
@@
The calculus contains a kinematical Bianchi identity. For details, see [Migdal(2025b)], Sec. 3.11, Eq. (43).
@@
The dot derivatives combine to form the loop operator $\mathcal{L}_{C}$, yielding a closed, linear diffusion equation for the Wilson loop functional. For details, see [Migdal(2025b)], Sec. 3.12, Eqs. (45-48).
@@ Appendix C Velocity Correlations and the Spectrum of Exponents
Here, $\zeta(s)$ is the Riemann zeta function and $f(z)$ is a calculable entire function defined in Appendix K in [Migdal(2024h)]. This function involves other functions $A(\Delta),B(\Delta),C(\Delta)$ that are defined in previous Appendices to that paper. These functions were precomputed in Mathematica ® and stored as fifth order interpolation through the dense table of results of high-precision numerical integration using Mathematica ® .
@@
Complete details of analytical and numerical computations involved in this large project are summarized in collection of Mathematica ® notebooks [Migdal(2024a), Migdal(2024b), Migdal(2024c), Migdal(2024d), Migdal(2024e), Migdal(2024f), Migdal(2024g)], which can be freely downloaded to verify these computations.
Appendix D The Hodge-Dual Matrix Surface
This section summarizes the definition of the Hodge-dual matrix-valued minimal surface. The full derivations and properties are detailed in Sec. 4.2 of [Migdal(2025b)].
Parametric Definition
The surface is embedded in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ with matrix-valued coordinates $X_{\mu}^{A}(\xi_{1},\xi_{2})$. For details, see [Migdal(2025b)], Sec. 4.2, Eq. (82).
$X_{\mu}=\tau_{A}X^{A}_{\mu}(\xi_{1},\xi_{2})=X^{0}_{\mu}(\xi_{1},\xi_{2})+\tau_{a}X^{a}_{\mu}(\xi_{1},\xi_{2})$ (100)
Scalar Area and Hodge-Duality
The area functional is defined with a Hodge-duality constraint. For details, see [Migdal(2025b)], Sec. 4.2, Eqs. (83-86).
$d\Sigma_{\mu\nu}$ $=d\xi_{1}d\xi_{2}\Sigma_{\mu\nu};$ (101)
$\Sigma_{\mu\nu}$ $=e_{lm}\partial_{l}X^{A}_{\mu}\partial_{m}X^{A}_{\nu};$ (102)
$|S|$ $=\int_{\mathcal{D}}\sqrt{\frac{1}{2}(\Sigma_{\mu\nu})^{2}};$ (103)
$\Sigma_{\mu\nu}$ $=\frac{1}{2}e_{\mu\nu\alpha\beta}\Sigma_{\alpha\beta};$ (104)
Dirichlet Boundary Conditions
The surface boundary is mapped from the loop $C(\theta)$ using ’t Hooft projectors. For details, see [Migdal(2025b)], Sec. 4.2, Eqs. (91-97).
$X^{A}_{\mu}(1,\theta)=\Lambda^{A}_{\mu\nu}C_{\mu}(\theta);$ (105)
$\Lambda^{0}_{\mu\nu}=\delta_{\mu\nu};\quad\Lambda^{a}_{\mu\nu}=\bar{\eta}^{a}_{\mu\nu};$ (106)
$\eta^{i}_{\mu\nu}=(\delta_{4\nu}\delta_{i\mu}-\delta_{4\mu}\delta_{i\nu})+e_{i\mu\nu 4};$ (107)
$\bar{\eta}^{i}_{\mu\nu}=(\delta_{4\nu}\delta_{i\mu}-\delta_{4\mu}\delta_{i\nu})-e_{i\mu\nu 4};$ (108)
$\frac{1}{2}e_{\mu\nu\lambda\rho}\eta^{i}_{\lambda\rho}=\eta^{i}_{\mu\nu};\quad\frac{1}{2}e_{\mu\nu\lambda\rho}\bar{\eta}^{i}_{\lambda\rho}=-\bar{\eta}^{i}_{\mu\nu};$ (109)
Conformal Gauge Condition
The conformal gauge is fixed by a condition on the induced metric at the boundary. For details, see [Migdal(2025b)], Sec. 4.2, Eq. (107).
$g_{lm}|_{\partial D}=\left(\partial_{l}\bar{X}^{A}_{\mu}\,\partial_{m}X^{A}_{\mu}\right)_{\partial D}\propto\delta_{lm}.$ (110)
Appendix E The loop equation for the minimal surface and its solution
This section summarizes how the Hodge-dual surface provides an exact, non-singular solution to the fixed-point loop equation. The full derivation is in Sec. 4.3 of [Migdal(2025b)].
Variational Principle
The area functional is defined as the minimum of a constrained action. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (111-113).
$|S|$ $=\min_{\lambda,X}\int d^{2}\xi\left(\sqrt{\frac{1}{2}\hbox{tr}\,\Sigma\cdot\Sigma^{T}}+\lambda_{i}\bar{\eta}^{i}_{\mu\nu}\Sigma_{\mu\nu}\right);$ (111)
$t_{\mu\nu}(r,\theta)$ $=\frac{\Sigma_{\mu\nu}}{\sqrt{2\hbox{tr}\,\Sigma\cdot\Sigma^{T}}}$ (112)
Area Derivative
The variation of the area functional $\delta|S|$ yields the area derivative. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (121, 122).
$\delta|S|$ $=\int d\theta\delta C_{\alpha}(\theta)\dot{C}_{\beta}(\theta)2T_{\alpha\beta}(1,\theta);$ (113)
$\frac{\delta|S|}{\delta\sigma_{\alpha\beta}(\theta)}$ $=2T_{\alpha\beta}(1,\theta)=2(t_{\alpha\beta}+\lambda_{i}\bar{\eta}^{i}_{\alpha\beta});$ (114)
The Hodge-Duality Constraint
The Hodge-duality of the area derivative requires the Lagrange multipliers to vanish at the boundary. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (123, 124).
$\bar{\eta}^{a}_{\mu\nu}\,T_{\mu\nu}(1,\theta)$ $=0\quad\forall a=1,2,3;$ (115)
$\lambda_{i}(1,\theta)$ $=0;$ (116)
The Solution of the Loop Equation
The area derivative of the Hodge-dual surface is, by construction, proportional to the self-dual tensor $t_{\mu\nu}$. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (125).
$\frac{\delta}{\delta\dot{C}_{\left[\mu\right.}(\theta-)}\frac{\delta}{\delta\dot{C}_{\left.\nu\right]}(\theta+)}|S|=2t_{\mu\nu}(1,\theta);$ (117)
This identity, combined with the kinematical Bianchi identity (from Appendix A), proves that the loop operator annihilates the area functional. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (126).
$\partial_{\mu}\frac{\delta|S|}{\delta\sigma_{\mu\nu}(\theta)}\propto e_{\mu\nu\lambda\rho}\left(\partial_{\alpha}(\theta)\right)\frac{\delta}{\delta\dot{C}_{\lambda}(\theta-)}\frac{\delta}{\delta\dot{C}_{\rho}(\theta+)}|S|\equiv 0$ (118)
Therefore, the exponential of the Hodge-dual area is an exact, non-singular solution to the fixed-point loop equation. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (127).
$\mathcal{L}_{\nu}(t)\,\exp\left(-\kappa S[C]\right)=\partial_{\mu}\left(\frac{\delta S[C]}{\delta\sigma_{\mu\nu}}\exp\left(-\kappa S[C]\right)\right)=0$ (119)
Planar Loop minimal surface
The Hodge-dual minimal surface can be analytically found for an arbitrary non-intersecting planar loop. The main result is the following theorem. We present this theorem here, correcting the proof in [Migdal(2025b)], where there was a typo in the boundary conditions for $Y$.
Theorem 2.
The minimal dual area for any non-intersecting planar loop equals $2\sqrt{2}$ times the planar area inside this loop.
Proof.
Let us use the conformal invariance of the dual area functional (101) and map the unit circle onto the given planar bounding loop $\mathbb{S}_{1}\mapsto C$. After that, the minimal surface can use the coordinates $\xi=(x_{1},x_{2})$ inside that planar loop $C$ as internal coordinates of the surface. In the context of our surface embedded into $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$, this parametrization reduces to
$X^{A}_{l}=\Lambda^{A}_{lm}\,\xi_{m},\quad l,m=1,2,$ (120)
$X^{A}_{\dot{l}}=Y^{A}_{\dot{l}}(\xi),\quad\dot{l}=3,4,$ (121)
$\xi\in\mathcal{D},\quad C=\partial\mathcal{D},$ (122)
where $\Lambda^{A}_{\mu\nu}$ are defined in (105), and the new field $Y^{A}_{\dot{l}}(\xi)$ satisfies the boundary conditions:
[…]
The area functional is
$|S|=\int_{\mathcal{D}}d^{2}\xi\sqrt{\tfrac{1}{2}\Sigma_{lm}^{2}+\Sigma_{l\dot{l}}^{2}+\tfrac{1}{2}\Sigma_{\dot{l}\dot{m}}^{2}},$ (124)
$\Sigma_{lm}=e_{ab}\Lambda^{A}_{al}\Lambda^{A}_{bm},$ (125)
$\Sigma_{l\dot{l}}=e_{ab}\Lambda^{A}_{al}\,\partial_{b}Y^{A}_{\dot{l}},$ (126)
$\Sigma_{\dot{l}\dot{m}}=e_{ab}\,\partial_{a}Y^{A}_{\dot{l}}\,\partial_{b}Y^{A}_{\dot{m}}.$ (127)
With the Dirichlet boundary values (123) the constrained (self–dual) minimizer of the area functional (124) is the linear map
$X^{A}_{\mu}(\xi)=\Lambda^{A}_{\mu m}\,\xi_{m},\qquad\xi\in D.$ (128)
This embedding automatically satisfies both the Hodge–duality constraint and the corrected boundary conditions. Its area tensor is constant,
$\Sigma_{\mu\nu}=\sum_{A}\Lambda^{A}E(\Lambda^{A})^{\!T}=2\!\begin{pmatrix}0&1&0&0\\[-3.0pt] -1&0&0&0\\[-3.0pt] 0&0&0&1\\[-3.0pt] 0&0&-1&0\end{pmatrix}$, (129)
which is self–dual ($\Sigma=\Sigma$) and yields $\Sigma_{l\dot{l}}=0$. The Euler–Lagrange system is identically satisfied for linear $X^{A}_{\mu}(\xi)$. The corresponding area density is $\sqrt{\tfrac{1}{2}\Sigma_{\mu\nu}\Sigma_{\mu\nu}}=2\sqrt{2},$ so that
$|S|=\!\!\int_{D}\!d^{2}\xi\,2\sqrt{2}=2\sqrt{2}\,|D|.$ (130)
Hence the minimal dual area for any non–intersecting planar loop equals $2\sqrt{2}$ times the Euclidean area enclosed by the loop, and the linear map (128) provides the unique regular self–dual minimizer consistent with the boundary data. ∎
Area Inequalities for Confinement
For an arbitrary loop $C$, the self-dual surface area $|S|$ is bounded by the unconstrained minimal area in the embedding space. These bounds are crucial for confinement. The full proofs are provided in [Migdal(2025b)] (Sec. 4.8).
Lower Bound
Theorem 3.
The area of the constrained matrix surface (i.e., with the duality constraint) is not smaller than that of the minimal unconstrained matrix surface with the same bounding curve. This follows from the fact that the constrained configuration space is a subset of the unconstrained space. See [Migdal(2025b)], Sec. 4.8, Eq. (168).
$\inf_{X\in\mathbb{R}^{4}\otimes\mathbb{R}^{4},\,\Sigma=\pm*\Sigma}|S|\geq\inf_{X\in\mathbb{R}^{4}\otimes\mathbb{R}^{4}}|S|.$ (131)
Upper Bound
Theorem 4.
The area of the unconstrained matrix surface is positive but not larger than four times the minimal surface area in Euclidean space ($A_{4}[C]$) with the same bounding curve $C\subset\mathbb{R}^{4}$. The proof relies on the inequality $\left|\Sigma_{\mu\nu}\right|\leq\sum_{A}\sqrt{\det\lVert g^{A}_{lm}\rVert}$. This leads to the final bound: See [Migdal(2025b)], Sec. 4.8, Eq. (172).
$0<\min_{X\in\mathbb{R}^{4}\times\mathbb{R}^{4}}|S|\leq 4A_{4}[C].$ (132)
For a planar loop, the Hodge-dual area is $2\sqrt{2}A_{4}[C]\approx 2.83A_{4}[C]$, which respects this bound.
Parity and the Final Solution
A key conceptual issue, first pointed out by E. Witten, is that a single Hodge-dual surface with a fixed chirality ($\lambda=\pm 1$) would violate parity. The full analysis is in [Migdal(2025b)] (Sec. 4.9).
To restore parity and preserve the crucial factorization property for separated loops, the solution must be the exponential of the sum of both the self-dual and anti-self-dual areas. See [Migdal(2025b)], Sec. 4.9, Eq. (176).
$W[C]=\exp\left(-\kappa\,(\,|S^{+}[C]|+|S^{-}[C]|\,)\right).$ (133)
This form is parity-even and solves the fixed-point loop equation because the loop operator, $\partial_{\mu}\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}$, annihilates each term in the exponent independently via the Bianchi identity (from Appendix A).
Conclusion. This provides a parity-preserving, factorizable, and exact analytic solution for the fixed point of the Yang-Mills gradient flow.

### v3

@@ Acknowledgements
I am grateful for insightful discussions regarding this theory with participants of seminars and conferences at the International Centre for Theoretical Sciences (ICTS) in Bangalore (India), where this theory was first reported, the Kavli Institute for Theoretical Physics (KITP), the Perimeter Institute for Theoretical Physics, Rutgers University, the Institute for Advanced Study (IAS), the University of Oxford, the Isaac Newton Institute for Mathematical Sciences in Cambridge, the London Institute for Mathematical Sciences, Imperial College London, the CERN Theoretical Physics Department, the "Turbulence on the Banks of the Arno" workshop in Pisa (2025), the 2nd European Fluid Dynamics Conference (EFDC2) in Dublin (2025), and numerous Zoom seminars in the US and Europe. Conversations with Katepalli Sreenivasan concerning the physics of turbulence, as observed in direct numerical simulations and laboratory experiments, were invaluable in clarifying the role and significance of loop space diffusion.
@@ Appendix A Essentials of the loop space calculus
The full derivations and technical details can be found in our recent paper [migdal2025SQYMflow].
@@
The area derivative is defined by the non-abelian Stokes theorem. For a detailed derivation, see [migdal2025SQYMflow], Sec. 3.3.
@@ A.1 Operator representation of the Wilson loop
The loop functional can be represented as a trace of a path-ordered exponential of covariant derivative operator. This subsection proves the fundamental operator identity that enables this. The full derivation, originally presented in [migdal2025SQYMflow] (Sec. 3.4), is reproduced here with all necessary mathematical details. We also fix some typos in that paper.
@@
The proof relies on discretizing the path and using the "disentangling" identity, which is a consequence of Feynman’s operator calculus [Feynman1951].
@@ 4. The Closed Loop Condition
The field strength $F_{\mu\nu}$ is generated by the antisymmetric part of the second dot derivative. For details, see [migdal2025SQYMflow], Sec. 3.6, Eq. (25).
@@
The covariant derivative of the field strength is generated by the discontinuity of the dot derivative, $\partial_{\mu}(t)\equiv\frac{\delta}{\delta\dot{C}_{\mu}(t-)}-\frac{\delta}{\delta\dot{C}_{\mu}(t+)}$. For details, see [migdal2025SQYMflow], Sec. 3.7, Eq. (31).
@@
The dot derivative discontinuity $\partial_{\mu}(t)$ annihilates the Wilson loop. For details, see [migdal2025SQYMflow], Sec. 3.9, Eq. (33).
@@
The area derivative acts as a first-order differential operator, satisfying the standard Leibniz rules. For details, see [migdal2025SQYMflow], Sec. 3.10, Eqs. (39-41).
@@
The calculus contains a kinematical Bianchi identity. For details, see [migdal2025SQYMflow], Sec. 3.11, Eq. (43).
@@
The dot derivatives combine to form the loop operator $\mathcal{L}_{C}$, yielding a closed, linear diffusion equation for the Wilson loop functional. For details, see [migdal2025SQYMflow], Sec. 3.12, Eqs. (45-48).
@@ Appendix C Velocity Correlations and the Spectrum of Exponents
Here, $\zeta(s)$ is the Riemann zeta function and $f(z)$ is a calculable entire function defined in Appendix K in [migdal2024quantum]. This function involves other functions $A(\Delta),B(\Delta),C(\Delta)$ that are defined in previous Appendices to that paper. These functions were precomputed in Mathematica ® and stored as fifth order interpolation through the dense table of results of high-precision numerical integration using Mathematica ® .
@@
Complete details of analytical and numerical computations involved in this large project are summarized in collection of Mathematica ® notebooks [MB40, MB41, MB42, MB43, MB44, MB45, MB46], which can be freely downloaded to verify these computations.